a car travel from one town to another at a speed 32 mph. if it had gone 4mph faster, it could have made the trip in 1/2 hr less time. how far apart are the town
Answer by mananth(16949)

a car travel from one town to another at a speed 32 mph. if it had gone 4mph faster, it could have made the trip in 1/2 hr less time. how far apart are the town
Let the distance between the towns be x
At 32 mph it takes x/32 hours. ( distance / speed)
If it had gone 4mph faster means it will go at 36 mph.
the time taken now will be x/36 hours
time taken is 1/2 hour less than earlier.
x/32-x/36= 1/2
36x-32x / 32*36= 1/2
4x= 32*36*1/2
4x= 576
x=576/4
x= 144 miles